Given Q=100K^0.5L^0.5 ,w= 30, r=40

i) Find the quantity of labour and capital that the firm should use in order to minimize the cost of

producing 1444 units of output

ii) What is this minimum cost?


1
Expert's answer
2021-10-03T14:14:36-0400

Given:

"Q = 100K^{0.5}L^{0.5}\\w =30\\r =4"

"MP_L=\frac{dQ}{dL}\\MP_L=100K^{0.5} × 0.5L^{-0.5}\\MP_L=50K^{0.5}L^{-0.5}\\MP_K = \frac{dQ}{dK}\\MP_K=100 × 0.5 K^{-0.5} L^{0.5}\\MP_K = 50K^{-0.5}L^{0.5}"

Cost minimizes at, 

"\frac{w}{r} = \frac{MP_L}{MP_K}\\\frac{30}{40}=\frac{50K^{0.5}L^{-0.5}}{50K^{-0.5}L^{0.5}}\\\frac{3}{4}=\frac{K}{L}\\3L=4K\\or,\\ K = \frac{3}{4}L\\L=\frac{4}{3}K\\K=\frac{3}{4}L"

Insert values in:

"Q = 100K^{0.5}×\frac{4}{3}K^{0.5}\\Q = \frac{400}{3} K\\When\space Q = 1444\\1444 = \frac{400}{3}K\\K =10.83\\ for L,\\Q = 100L^{0.5}×\frac{3}{4}L^{0.5}\\Q = \frac{300}{4} L\\When\space Q = 1444\\ 1444=\frac{300}{4} L\\L =19.25"


b). 

"C = wL + rK\\

C =30×19.25+40×10.83\\

C = 1010.7"